Fermi National Laboratory


Accelerator Update for November 30, 2005 - December 2, 2005

Wednesday November 30

The day shift began with the Tevatron (TeV) sector B17 work continuing, with Operations monitoring a very small Antiproton Source (PBR) stack, with the Recycler conducting studies, and with MiniBooNE, NuMI, and the Meson MTest and MCenter experiments taking beam.

Pbar experts began a reverse proton study at 8:16 AM.

At 12:26 PM, Operations reported a Booster LCW pressure drop. Investigations showed that a pump at CUB had tripped off. Operators had already turned much of Booster off, due to the pressure drop, but, after the pump came back online, began turning Booster on around 12:50 PM. Booster was back to normal operations by 1:05 PM.

At 1:05 PM, a Linac RF station (LRF3) tripped off. Operators had it back on ten minutes later.

At 1:26 PM, LRF3 tripped off again.

At 1:31 PM, Operations reported that a Booster RF station (BRF16) had tripped off four times in the last ten minutes. It reset okay.

Operations reported at 2:04 PM that the TeV sector F0 work was done and the P1 transfer line power supplies were up and running.

The Main Injector (MI) began a 2.5 Mhz study at 2:35 PM.

At 2:50 PM, BRF16 tripped off. It reset okay.

At 6:57 PM, BRF10 tripped off for the first of five times on a ground fault indication. It reset each time.

Recycler experts completed their studies at 10:33 PM.

At 11:40 PM, Operations reported that they had bypassed BRF10 due to all its ground fault trips.

Thursday December 1

The midnight shift began with TeV B17 work continuing, with Pbar conducting reverse proton studies, with the Recycler conducting Electron Cooling studies, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

The MTest experiment had their beam shut off at 8:28 AM.

At 11:09 AM, Operations reported losing almost all of the MI RF stations. A RF expert reported that reflective power had caused the trip.

MTest requested beam at 11:44 AM.

At 2:32 PM, BRF10 tripped off and wouldn’t reset. Operators bypassed the station and brought BRF19 online.

At 3:42 PM, LRF2 tripped off and wouldn’t reset. Operations investigated.

An MI expert began a 2.5 MHz study at 6:06 PM.

A TeV expert reported at 8:47 PM that he had found a bad feed-through on a B11 separator. Earlier, a vacuum leak was found on the B13 half-cell.

Friday December 2

The midnight shift began with TeV B17 work continuing, with Pbar conducting reverse proton studies, with the Recycler conducting studies, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

Cryo system technicians began a TeV sector B2 wet engine overhaul at 6:51 AM.

The Plans for Friday and the Weekend

The plans are to continue TeV sector B1 work while delivering beam to the other experiments. Run Coordinator reports that after the separator feed-through is installed, the separator will require a bake-out, which will run until December 9, and then there must be a conditioning period. He estimates that beam should return on December 12.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/2/2005   email Fermilab